OpenStudy (anonymous):

What is the square root of 81

OpenStudy (aripotta):

use a calculator, though i think everyone should know this

hero (hero):

What's 9 x 9 = ?

OpenStudy (anonymous):

well gezz i thought thats what the question bored was for.. questions

OpenStudy (aripotta):

on the calculator, just put in 81, then push the square root button

OpenStudy (aripotta):

well, i'm just saying...sorry if that sounded rude or anything, i didn't mean to be.

OpenStudy (anonymous):

were cool :D

OpenStudy (aripotta):

i'm guessing you have the answer now? or do you still need help?

OpenStudy (anonymous):

no thank you i have it thanks for the help

OpenStudy (wwe123):

the square root of 81 is +9 or -9

OpenStudy (wwe123):

+9 and -9

OpenStudy (aripotta):

\[\pm9\]

OpenStudy (aripotta):

:)

Directrix (directrix):

@wwe123 @AriPotta @grapehead The square root of 81 refers to the principal square root (positive) of 81 which is 9 and only 9. If you are solving the equation x^2 = 81, then x = 9 or x = -9.

OpenStudy (aripotta):

i'm sure it could be either/or. it's technically correct.

hartnn (hartnn):

directrix is correct

hartnn (hartnn):

x=\(\sqrt{81}\) can only have 1 root because its the equation with degree=1

hartnn (hartnn):

when you square it, u attach an extra root to it

hartnn (hartnn):

x^2=81 had 2 roots

OpenStudy (aripotta):

......ok

OpenStudy (aripotta):

-_- i got that. thanks.

OpenStudy (wwe123):

no 9*9 =81 -9*-9=81 so, \[\sqrt{81} =\pm9\]

hartnn (hartnn):

@wwe123 please clear your basic algebra concepts.

OpenStudy (wwe123):

@Directrix

OpenStudy (aripotta):

omg just drop it.

OpenStudy (aripotta):

and you guys are freaking deleting your posts again. that really aggravates me

hartnn (hartnn):

@AriPotta we just want u to have clear basics so that u do not misguide others while helping....

OpenStudy (aripotta):

yea, and i said i understood it.

OpenStudy (aripotta):

so idk why we keep discussing it. and why you guys are deleting comments.